# EquiReg: Equivariance Regularized Diffusion for Inverse Problems

### v1

@@ 3 Preliminaries
The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the actions in the group $G$ is equivariant if $S_{g}(f({\bm{z}}))=f(T_{g}({\bm{z}}))$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, and $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
While prior work leverages equivariance to directly incorporate symmetries into deep neural networks [90, 91, 92, 94, 112, 113], recent studies explore approximate equivariant networks to relax strict mathematical symmetries that may not fully hold in real-world data, aiming to improve performance [114]Wang et al. [114] proposes a definition of approximate equivariance (Definition 3.2[114], which reduces to strict equivariance when $\epsilon=0$ (Definition 3.1). For such functions, we define the equivariance error (Definition 3.3) to quantify the deviation from perfect symmetry.
Definition 3.2 (Approximate Equivariance).
Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$. The function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\epsilon$-approximate equivariant if $\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|\leq\epsilon$ for all ${\bm{z}}\in\mathcal{Z}$ and $g\in G$ where $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$.
Definition 3.3 (Equivariance Error).
The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ is $\mathcal{L}_{\text{EE}}(f)=\sup_{{\bm{z}},g}\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating on the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ for $g\in G$. Hence, $f$ is $\epsilon$-approximate equivariant iff $\mathcal{L}_{\text{EE}}(f)<\epsilon$.
Data manifold hypothesis. Let data ${\bm{x}}\in\mathcal{X}\subset\mathbb{R}^{d}$ be in an ambient space of dimension $d$ with support $\mathcal{X}$ distribution. We assume that data are sampled from a low-dimensional manifold $\mathcal{M}$ [115, 116] embedded in a high-dimensional space (Assumption 3.1). This data hypothesis is popular in machine learning [117], and has been studied mathematically in the literature [118, 116, 119, 120]. Moreover, empirical evidence in image processing supports the manifold hypothesis [121, 122], and prior diffusion-based inverse solvers assume this manifold property [62, 33, 32].
Assumption 3.1 (Manifold Hypothesis).
@@ 4 EquiReg: Equivariance Regularized Diffusion
We first discuss our proposed generalized regularization framework for improving diffusion-based inverse solvers. Then, we introduce a new class of functions with distribution-dependent equivariance errors. Lastly, we propose two objective losses to regularize diffusion models for inverse problems.
Generalized regularization of diffusion models for inverse problems. EquiReg addresses the propagation error caused by the posterior approximation $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ using regularization. This regularization can be used as a plug-in on a diverse range of pixel-based and latent diffusion models. We explain EquiReg in the context of Wasserstein Gradient Flows.
Proposition 4.1.
@@
Because the conditional expectation $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{T-t}]$ is a linear combination of all candidate ${\bm{x}}_{0}$, the approximation remains relatively accurate when $T-t$ is small (i.e., ${\bm{x}}_{T-t}$ stays close to the data manifold under low noise) but may incur high error for larger $T-t$, as shown in Figure 3. To mitigate this, we reweight the contributions to the first term of $\Phi$, down-weighting unreliable estimates, and amplifying the reliable ones. The resulting reweighted functional is
@@
Proposition 4.2.
(Informal) The evolution of $\rho$, the probability distribution of ${\bm{x}}_{T-t}$ driven by the practical and regularized reverse dynamics (eq. 5), is an approximation of the Wasserstein-2 gradient flow associated with minimizing $\tilde{\Phi}$.
@@
In contrast to our work, most prior approaches focus on locally reducing likelihood error and do not adopt a generalized view of correcting the full functional being minimized. Next, we introduce distribution-dependent equivariance error and show how to design a regularizer with low error on or near the manifold and high error off-manifold.
Distribution-dependent equivariance errors. While the literature has primarily studied the equivariance properties of functions for general inputs, we propose a new equivariance error for approximately equivariant functions under the support of an input data distribution (see Definition 4.1).
Definition 4.1 (Distribution-Dependent Equivariance Error).
The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the distribution $p$ is $\mathcal{L}_{\text{EE}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|]$ where $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ is the set of transformations under the symmetry group $G$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$, for $g\in G$.
The above definition enables us to define the equivariance error of trained function mappings for both on and off the manifold data. We leverage this definition in the next section to find functions whose equivariance error is low for on-manifold data and high elsewhere. Moreover, we introduce a new constrained distribution-dependent equivariance error, where the input is implicitly regularized to lie on the manifold $\mathcal{M}$ in addition to minimizing the equivariance error (Definition 4.2). Both equivariance errors are non-local, defined at the distribution level. When used to regularize the reverse conditional diffusion process, they are computed via local evaluations over the sampled data.
Definition 4.2 (Manifold Constrained Distribution-Dependent Equivariance Error).
Let $T_{g}:\mathcal{Z}\rightarrow\mathcal{Z}$ define the set of transformations under the symmetry group $G$, where $g\in G$. The manifold constrained equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the group $G$ and under the data distribution $p$ is $\mathcal{L}_{\text{EE}}^{\mathcal{M}}(f,p)=\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|{\bm{z}}-h(S_{g}^{-1}(f(T_{g}({\bm{z}}))))\|]$ where $h:\mathcal{X}\rightarrow\mathcal{Z}$, $S_{g}$ is the transformation corresponding to $T_{g}$ operating the output space $S_{g}:\mathcal{X}\rightarrow\mathcal{X}$ of function $f$, and decoder-encoder composition of $f$-$h$ forms a vanishing-error autoencoder under the manifold $\mathcal{M}$ (see Appendix for the definition of vanishing-error autoencoders and their usage in prior work).
(a) Training induced equivariance.
(b) Data inherent equivariance.
Figure 4: MPE function examples.
Proposed regularization. Given the aforementioned formulated errors, we are now ready to discuss the choice the function $f$. We term a class of manifold-preferential equivariant (MPE) functions, whose equivariance error is lower for samples on the data manifold than for off-manifold samples. We consider two approaches for identifying such functions.
Training induced equivariance In this case, MPE emerges via symmetry-preserving augmentations applied during training. Prior work has studied equivariant properties of learned representations in deep networks [123], showing that data augmentations [124] and representation compression by reducing the model capacity [125] promote equivariant features even when equivariance is not explicitly built into the architecture. Importantly, the trained network is only approximately equivariant, and prior studies have sparsely noted that symmetry-preserving properties degrade for inputs deviating from in-distribution data [126]. A few studies have leveraged this emergence of MPE on trained networks for out-of-distribution detection [127, 128, 129].
Figure 4(a) illustrates the MPE property, emergent via training induced equivariance, of $\mathcal{E}$-$\mathcal{D}$ of a pre-trained variational autoencoder, currently used in LDMs [61]. Specifically, it shows that the equivariance error is lower for natural images and increases when images deviate from the clean data distribution. Based on Definitions 4.1 and 4.2, we propose Equi and EquiCon losses using a pre-trained encoder-decoder for diffusion-based inverse solvers such as DPS, PSLD, and ReSample.
$\text{Equi}_{\text{pixel}}\quad\mathcal{R}({\bm{x}}_{t})$ $=\|S_{g}(\mathcal{E}({\bm{x}}_{0|t}))-\mathcal{E}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$ (6)
$\text{Equi}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|S_{g}(\mathcal{D}({\bm{z}}_{0|t}))-\mathcal{D}(T_{g}({\bm{z}}_{0|t}))\|_{2}^{2}$
$\text{EquiCon}_{\text{latent}}\quad\mathcal{R}({\bm{z}}_{t})$ $=\|{\bm{z}}_{0|t}-\mathcal{E}(S_{g}^{-1}(\mathcal{D}(T_{g}({\bm{z}}_{0|t}))))\|_{2}^{2}$
Algorithm 1 Equi-DPS for Image Restoration Tasks
@@
2:${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
3:for $t=T-1$ to 00 do
4:    $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
5:    ${{\bm{x}}}_{0|t}\leftarrow\frac{1}{\sqrt{\bar{\alpha_{t}}}}({\bm{x}}_{t}+(1-\bar{\alpha_{t}})\hat{\bm{s}})$
@@
We use the above regularization for image restoration tasks (e.g., see Algorithm 1 for Equi-DPS).
Data inherent equivariance MPE can also emerge naturally from symmetries underlying the data. This often occurs in physics systems where the combination of coefficient functions, boundary values, and solution functions of PDEs remains valid after applying an invertible coordinate transformation consistently to them. Let $\mathcal{G}(a)\mapsto u$ be a PDE operator that maps from the initial condition $a$ to the solution $u$. Then, it follows that $S_{g}(\mathcal{G}(a))=\mathcal{G}(T_{g}(a)).$

### v2

@@ 2 Preliminaries and Related Works
Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$. A function $f:\mathcal{Z}\to\mathcal{X}$ is equivariant if for all $g\in G$ and ${\bm{z}}\in\mathcal{Z}$, $f(T_{g}({\bm{z}}))=S_{g}(f({\bm{z}}))$.
An equivariant function preserves structure under group transformations (Definition˜2.1). While prior work leverages exact equivariance to encode symmetries directly into neural networks, recent studies investigate approximate equivariance to relax strict symmetry assumptions that may not hold in real-world data, aiming to improve performance (Wang et al., 2022). These works introduce a formal definition of approximate equivariance (Definition˜2.2) and an equivariance error to quantify deviations from perfect symmetry.
Definition 2.2 (Approximate Equivariant Functions).
Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$. A function $f:\mathcal{Z}\rightarrow\mathcal{X}$ is $\epsilon$-approximate equivariant if for all $g\in G$ and ${\bm{z}}\in\mathcal{Z}$, $\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|\leq\epsilon$. The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ is defined as $\sup_{{\bm{z}},g}\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|$. Hence, $f$ is $\epsilon$-approximate equivariant iff its error $\epsilon$.
We use the term manifold which refers to the data manifold hypothesis (see Assumption˜H.1(Cayton and others, 2005) that assumes data is sampled from a low-dimensional manifold embedded in a high-dimensional space. This hypothesis is popular in machine learning (Bordt et al., 2023) and diffusion-based solvers (He et al., 2024; Chung et al., 2022b; 2023), supported by empirical evidence for imaging (Weinberger and Saul, 2006).
@@ 3 EquiReg: Equivariance Regularized Diffusion
We begin by presenting a generalized regularization framework for improving diffusion-based inverse solvers. We focus on the property of equivariance and introduce a new class of functions whose equivariance errors are distribution-dependent (low for on- or near-manifold samples and high for off-manifold samples). We leverage these functions to regularize diffusion models, guiding sampling trajectories toward better inverse solutions.
This paper addresses the propagation error introduced by the approximation of posterior $p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ by incorporating an explicit regularization term. The proposed framework is general and can be applied as plug-in on a wide range of pixel and latent-space diffusion models. Given $p_{t}({\bm{y}}|{\bm{x}}_{t})=\int{p({\bm{y}}|{\bm{x}}_{0})p_{t}({\bm{x}}_{0}|{\bm{x}}_{t})\mathrm{d}{\bm{x}}_{0}}$, let $\tilde{p}_{t}({\bm{x}}_{0}|{\bm{x}}_{t})$ denote an approximation of the posterior to make the likelihood tractable. We formulate the regularized reverse diffusion dynamics as
@@
Algorithm 1 Equi-DPS for Inverse Problems.
@@
3:for t=T1t=T-1 to 0 do
4:  $\hat{\bm{s}}\leftarrow\bm{s}_{\theta}({\bm{x}}_{t},t)$
@@
where $\mathcal{R}({\bm{x}}_{t})$ is the regularizer. Applying this to DPS (Chung et al., 2023) takes the form in Algorithm˜1). This formulation brings us to the key contribution: how to design the regularizer to be effective in improving posterior sampling of diffusion models. Regularization in optimization aims to penalize undesired solutions, moving the algorithm towards solutions with low regularizer value. In the context of diffusion models for sampling from data distributions, we interpret an ideal regularizer as follows: it should yield low values for on-manifold and high values for off-manifold samples, enabling accurate posterior sampling even when the likelihood score is approximated.
In terms of sampling dynamics, i.e., when applied at each reverse-diffusion step, the regularizer should effectively penalize trajectories leaving the data manifold and reinforce those aligned with high-probability regions. This motivates designing a regularizer that applies a global correction to the entire functional, in contrast to prior works that focus only on locally reducing likelihood error. The ideal property of a regularizer would be to produce high error on undesirable samples and low error on desirable samples. We use equivariance, a global property that enforces geometric symmetries to instantiate such a regularizer to guide the diffusion process toward the data manifold. To realize this idea, we seek functions that exhibit approximate equivariance and discriminate on- from off-manifold samples.
We propose to quantify the equivariance of a function relative to a data distribution. Specifically, while the literature has primarily studied the equivariance properties of functions for general inputs, we propose a new definition for functions in which their equivariance error is distribution-dependent and defined under the support of an input data distribution (Definition˜3.1).
Definition 3.1 (Distribution-Dependent Equivariant Functions).
Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$. The equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the distribution $p$ is defined as $\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|S_{g}(f({\bm{z}}))-f(T_{g}({\bm{z}}))\|]$.
The above definition enables us to define functions whose equivariance error can differentiate on-manifold samples from off-manifold ones. We aim to find functions whose equivariance error is low for on-manifold data and high elsewhere. We also introduce a constrained version of equivariance error, where the input is implicitly regularized to lie on the manifold $\mathcal{M}$ in addition to minimizing the equivariance error (Definition˜3.2). Both equivariance errors are non-local, defined at the distribution level. When used to regularize the reverse conditional diffusion process, they are computed via local evaluations over the sampled data.
Definition 3.2 (Manifold-Constrained Distribution-Dependent Equivariant Functions).
Let $G$ act on $\mathcal{Z}$ via $T_{g}:\mathcal{Z}\to\mathcal{Z}$ and on $\mathcal{X}$ via $S_{g}:\mathcal{X}\to\mathcal{X}$. The manifold-constrained equivariance error of the function $f:\mathcal{Z}\rightarrow\mathcal{X}$ under the data distribution $p$ is $\sup_{g}\mathbb{E}_{{\bm{z}}\sim p}[\|{\bm{z}}-h(S_{g}^{-1}(f(T_{g}({\bm{z}}))))\|]$ where $h:\mathcal{X}\rightarrow\mathcal{Z}$, and the pair $(f,h)$ forms a vanishing-error autoencoder (see Appendix I).
To define our method, we term a class of manifold-preferential equivariant (MPE) functions, whose equivariance error is lower for samples on the data manifold than for off-manifold samples. EquiReg is a regularization framework, not a manifold projection method. EquiReg penalizes states that deviate from symmetry-preserving regions; when an MPE function is used, these regions align with the data manifold. In practice, MPE functions can emerge in different ways, which we illustrate with examples from augmented training and from data symmetries. MPE can emerge when functions are trained with symmetry-preserving mechanisms such as data augmentation. Prior work has studied equivariant properties of learned representations in deep networks (Lenc and Vedaldi, 2015), showing that data augmentations (Krizhevsky et al., 2012) and representation compression via reduced model capacity (Bruintjes et al., 2023) promote equivariant features even when equivariance is not explicitly built into the architecture. Importantly, the trained network is only approximately equivariant, and prior studies have noted that symmetry-preserving properties degrade for inputs deviating from in-distribution data (Azulay and Weiss, 2019). A few studies have leveraged this emergent MPE in trained networks for out-of-distribution detection (Zhou, 2022; Kaur et al., 2022; 2023).
To demonstrate the widespread MPE property of learned mappings, we have considered additional pre-trained models and quantified their equivariance loss for several datasets, i.e., natural images and corrupted ones (see Appendix˜I of Appendix.) Figure˜4(a) illustrates the MPE property, emergent via training with augmentations, of $\mathcal{E}$-$\mathcal{D}$ of a pre-trained autoencoder, currently used in LDMs. Specifically, it shows that the equivariance error is lower for natural images and increases when images deviate from the clean data distribution. Based on Definitions˜3.1 and 3.2, we propose Equi loss using an MPE function $f$ for diffusion-based inverse solvers:
$\mathcal{R}_{f_{\text{MPE}}}({\bm{x}}_{t})$ $=\|S_{g}(f_{\text{MPE}}({\bm{x}}_{0|t}))-f_{\text{MPE}}(T_{g}({\bm{x}}_{0|t}))\|_{2}^{2}$ (5)
where ${\bm{x}}_{0|t}$ and ${\bm{z}}_{0|t}$ are function of ${\bm{x}}_{t}$ and ${\bm{z}}_{t}$, respectively. MPE can also emerge due to symmetries present in the data itself during training. This often occurs in physics systems where coefficient functions, boundary values, and solution functions of PDEs remain valid under invertible coordinate transformations. Formally, let $\mathcal{G}(a)\mapsto u$ be a PDE operator mapping initial condition $a$ to solution $u$, and let $T_{g}$ and $S_{g}$ be invertible transformations that preserve PDE structure and boundary conditions. Then, $S_{g}(\mathcal{G}(a))=\mathcal{G}(T_{g}(a))$. Neural operators (Kovachki et al., 2021), popular architectures for modelling physics, trained on PDEs with such inherent symmetries can learn equivariance properties. Figure˜4(b) shows that an MPE function can be constructed using Fourier Neural Operators (FNOs (Li et al., 2021)) trained on non-augmented Navier–Stokes physics data. When a pre-trained FNO is used as the MPE function, the equivariance loss in equation 5 is lower for in-distribution data than for out-of-distribution data under reflection as the group action.
(a) Equivariance from augmentation, error computed with $f_{\text{MPE}}=\mathcal{E}$ from autoencoder pretrained on FFHQ-256 and evaluated on examples from FFHQ-256.
(b) Equivariance arising from data symmetries. Error computed with $f_{\text{MPE}}=\text{FNO}$ and evaluated on Navier-Stokes data.
Figure 4: Equivariance error is consistently lower for clean vs perturbed examples.
The key message from our MPE examples is that MPE behaviour naturally emerges when a function (e.g., a neural network) is trained with appropriate augmentations or when the data exhibit inherent symmetries. Our results extend this analysis and demonstrate that MPE functions are widely present and can be effectively integrated into the EquiReg framework to improve posterior sampling (Table˜8). We leverage this property to distinguish on-manifold from off-manifold samples and to regularize the posterior sampling trajectory toward high-probability regions. Finally, we note that the choice of symmetry group may often be a challenge depending on application domain, a shared challenge in the broader equivariance literature. We provide guidelines on how to choose symmetry groups in Appendix˜H with reference on automatic symmetry discovery from data (Zhou et al., 2021; Quessard et al., 2020; Dehmamy et al., 2021; Mohapatra et al., 2025).
Finally, we complement our empirical results with a theoretical discussion of the role of the function $\mathcal{R}$ in regularizing conditional diffusion models and its impact on the sampling trajectory. In particular, we gain insight into the desirable properties of an optimal regularizer by reinterpreting the reverse conditional diffusion process as a time-inhomogeneous Wasserstein gradient flow (Ferreira and Valencia-Guevara, 2018) (see Propositions˜G.1 and G.2 in Appendix).
Table 1: EquiReq Characterization under change of EquiReg period, $\lambda_{t}$, and DDIM steps.
Super Resolution Gaussian Blur
[…]
DPS N/A 46.20 22.99 (1.93) 0.20 (0.05) 135.7 46.50 24.59 (2.25) 0.15 (0.03) 88.70
Equi-DPS 1 51.10 26.73 (1.99) 0.12 (0.03) 87.97 52.20 26.08 (2.25) 0.12 (0.03) 87.11
Equi-DPS 2 48.90 26.73 (1.99) 0.12 (0.03) 87.98 49.10 26.06 (2.24) 0.12 (0.03) 87.19
Equi-DPS 5 47.10 26.73 (1.99) 0.12 (0.03) 87.98 47.30 26.06 (2.24) 0.12 (0.03) 87.32
Equi-DPS 10 46.90 26.73 (1.99) 0.12 (0.03) 87.99 47.00 26.05 (2.24) 0.12 (0.03) 87.04
(a) Robustness and computational efficiency of applying EquiReg under various periods during sampling. EquiReg maintains performance when applied every $\{1,2,5,10\}$ DDIM steps while incurring minimal computational overhead.
DPS PSLD
[…]
0.0 24.34 (1.03) 0.664 (0.061) 0.156 (0.051) 0.0 23.83 (2.61) 0.63 (0.12) 0.315 (0.07)
0.001 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.01 25.35 (2.24) 0.70 (0.09) 0.280 (0.07)
0.01 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.1 26.63 (1.68) 0.74 (0.08) 0.337 (0.06)
0.1 25.44 (1.22) 0.708 (0.057) 0.118 (0.038) 0.25 26.22 (1.57) 0.72 (0.08) 0.366 (0.05)
1.0 25.44 (1.22) 0.709 (0.057) 0.118 (0.038) 1.0 24.74 (1.28) 0.66 (0.07) 0.438 (0.05)
(b) Robustness of EquiReg to the choice of $\lambda_{t}$. Sensitivity analysis for DPS and PSLD.
DPS Equi-DPS (ours)
[…]
500 13.89 0.0937 0.955 417.07 20.61 0.366 0.500 238.51
750 21.77 0.540 0.254 153.74 25.60 0.704 0.160 110.89
900 22.97 0.628 0.201 148.03 26.52 0.755 0.126 88.46
1000 22.99 0.649 0.201 135.71 26.73 0.767 0.120 88.00
(c) EquiReg improves performance under reduced DDIM steps. Pixel-based super-resolution on FFHQ $256\times 256$.
Table 2: EquiReg improves SITCOM under reduced measurement consistency steps ($K_{\text{meas}}$). We reduce $K_{\text{meas}}$ and add an equal amount of EquiReg steps ($K_{\text{EquiReg}}$). Evaluated with $50$ DDIM steps on motion deblur for FFHQ.
[…]
10 N/A 28.06 0.81 21.57
5 5 29.26 0.83 11.09
20 N/A 27.04 0.79 38.85
10 10 28.93 0.82 20.92
30 N/A 27.79 0.80 58.84
15 15 29.63 0.84 30.19
40 N/A 30.40 0.85 78.08
20 20 29.50 0.83 41.02
60 N/A 28.35 0.81 108.57
30 30 31.36 0.87 59.38
4 Results
This section provides experimental results evaluating the performance of EquiReg on inverse problems, including linear and nonlinear image restorations and PDE solving. To fairly assess the impact of EquiReg, we adopt a paired comparison setting (e.g., PSLD vs. Equi-PSLD) across experiments, keeping all other factors such as architecture, training, and sampling fixed. This design ensures that any observed improvements can be attributed specifically to EquiReg rather than to differences in the underlying model or inference procedure. For fair comparison, we set a fixed seed for the initial noise and evaluate the impact of EquiReg under reduced measurement consistency and sampling steps, providing a path toward faster diffusion-based inversion. Results emphasize EquiReg’s usefulness when the baseline performance deteriorates. We analyze the diversity trade-offs, and demonstrate the broad emergence of MPEs and their usefulness within EquiReg. Lastly, we provide preliminary analysis on EquiReg improving the realism of text-guided image generation.
Image restoration tasks. We evaluate EquiReg when applied to: SITCOM (Alkhouri et al., 2025), PSLD (Rout et al., 2023), ReSample (Song et al., 2023a), and DPS (Chung et al., 2023). We compare against manifold-preserving or geometry-constraint approaches including MCG (Chung et al., 2022b), MPGD-AE (He et al., 2024), and DiffStateGrad (Zirvi et al., 2025).